# CleAR: Robust Context-Guided Generative Lighting Estimation for Mobile Augmented Reality

### v1

@@ 7. User Study
Our user study is conducted as an online survey via the Qualtrics [^7] Qualtrics: https://www.qualtrics.com/ platform. Our studies were approved by our organization’s Institutional Review Board (IRB) and then distributed through personal and professional networks. Participants in this study were voluntary and subject to informed consent, presented at the beginning of the survey form.
@@
Figure 14. Visual quality assessment user study results. We compare the virtual object rendering visual quality impacts between CleAR and four other methods from 31 participants’ feedback in six different environments. Overall, participants report the visual quality rating of CleAR ranked the top with an average score of 4.87, 12% higher than the second-best method, StyleLight (average score: 4.3). The overall standard deviation of the rating values for CleAR (1.56) is also lower than StyleLight (1.67) by 7%, showing more robust estimation quality.
7.1. Study Protocol
Our study survey consists of three main chapters. The first chapter surveys the participants’ past experiences with mobile AR and their impressions of virtual object rendering qualities with existing mobile AR applications. We received responses from 31 individuals (48% female, 45% male, 6% non-binary/third gender) with varying levels of expertise in computer technology. Respondents were familiar with multiple areas of graphics technology, with image editing being the most common (51.6%). This was followed by video editing (41.9%) and 3D modeling and rendering (35.5%). 25.8% of respondents reported not being familiar with specific areas of graphics technology. 58% individuals have used AR devices. Among the respondents, mobile phones or tablets were the most commonly used devices for AR applications, with 48% having used them. Meta Quest was the second most popular device, used by 29% of the respondents.
The second chapter of our survey is a training section that guides participants in completing the quality assessment study. This chapter first shows a quality rating question where the participants will be given 5 images. Each image contains the same virtual sphere rendered by five different lighting conditions created from five sources. Participants are instructed to rate the visual qualities using a Likert scale of 1 to 7, where 1 represents the lowest quality and 7 represents the highest. Participants are not aware of which lighting sources are used and their mappings to the images. An example rating for this question is shown to the user for training purposes. Next, we will show a follow-up question to ask participants for feedback, specifically on the quality issues of images generated by our system. Following the training chapter, the third is the formal quality assessment study.
7.2. Results
@@
In the formal study, we presented participants with six question groups, each with a distinctive virtual object. The question group follows the format used in the training chapter. The order of environment lighting used for generating the images is randomized and will not be shown to participants. In Figure 14, we show six groups of visual quality ratings. Each group uses different environments and virtual object setups, representing environments with varying lighting conditions and objects with various geometries and materials. Six virtual objects are used in total, including a diffuse sphere, matte sphere, mirror sphere, floor lamp, tea set, and coffee mug. Specifically, lighting conditions used in these questions were generated using CleAR, DiffusionLight (Phongthawee et al., 2023), StyleLight (Wang et al., 2022), OmniDreamer (Akimoto et al., 2022), and physically captured HDR environment maps.
@@
Excluding the ratings of virtual objects rendered by physically captured lighting, from the collected responses, 4 out of 6 objects reported better virtual object rendering quality using environment lighting generated by CleAR. On average, the visual quality rating of CleAR ranked the top with an average score of 4.87 across all responses. This score is about 12% higher than the second-best method, StyleLight (average score: 4.35). Additionally, the overall standard deviation of the rating values for CleAR (1.56) is also lower than StyleLight (1.67) by 7%, showing more robust estimation quality.
@@
Interestingly, we found that, in two questions (tea set and coffee mug), participants’ ratings suggest even better visual quality for CleAR than the physically captured environment maps. Upon further inspection, we found that, although physically captured environment lighting represents a more accurate of the radiance conditions in the scene, human vision favors the slightly brighter outputs generated by CleAR. We hope this finding raises new interest in the research community to design and evaluate lighting estimation systems in the future with considerations of human perception.
@@ 8.1. Environment Lighting and Rendering
Pioneering works have established several ways of capturing high-fidelity physical environment lighting and representing them in digital formats. Omnidirectional HDR environment map is a commonly adopted solution (Debevec, 2006) for representing the environment lighting because it stores the incoming radiance information that can be easily integrated with modern computer graphics rendering. Environment maps can be commonly captured using mirror balls, 360 cameras, or bracketed image stitching (Debevec and Malik, 2023; Brown and Lowe, 2007). A recent work, GLEAM (Prakash et al., 2019), incorporates the mirror ball-assisted environment lighting capturing process into AR applications. However, the requirement for the presence of the mirror ball limits the practicality of AR device usage. Instead, our system uses the generative estimation approach to provide flexible, high-quality lighting estimation. Additionally, HDR environment maps can be challenging to obtain on AR devices due to many sensor limitations. Recent lighting estimation systems (Zhao and Guo, 2021; Zhao et al., 2022) can only output LDR ones. But our two-stepped generative estimation design allows CleAR to support HDR lighting estimation.
8.2. Image Generative Models
@@
Recently, generative models have demonstrated impressive capabilities in image synthesis, as well as generating audio, 3D models, and other data modalities (Yang et al., 2023; Croitoru et al., 2023). The generative diffusion model (Ho et al., 2020) receives particularly strong attention from the research community because of its high-quality outputs in image generation. The generative diffusion model uses a physical diffusion process inspired by the Gaussian signal denoising generation process. Combining this novel generation process with the large model parameter sizes, generative diffusion models outperform generative models with prior architectures, such as GAN (Goodfellow et al., 2020) or VAE (Esser et al., 2021). Several recent works (Chen et al., 2022; Akimoto et al., 2022; Wang et al., 2022; Phongthawee et al., 2023) also propose to adopt generative models, including diffusion models, to solve the generic lighting estimation problem. In this work, we focus on solving the lighting estimation problem in the context of mobile AR applications using generative models.
8.3. Context-aware Mobile System
@@
Context-aware computing is a classical computing paradigm in which applications can discover and take advantage of contextual information (Chen and Kotz, 2000). Early research in context-aware AR systems (Starner et al., [n. d.]) demonstrated that important environment information can be extracted from camera frames for task planning and decision-making. In recent years, new developments of AR systems have also sought to leverage broad types of environment context information to assist several AR tasks and achieve better user experiences (Lam et al., 2021; Tahara et al., 2020; Wang et al., 2020). Our work uses two types of context information: environment semantics and ambient light sensor data. These context data provide important complementary information and important guidance for matching the environment map visual context and the lighting conditions between the estimated result and the physical lighting conditions.
9. Conclusions
@@
In this work, we introduced an end-to-end lighting estimation framework CleAR that can be easily integrated into many existing mobile AR applications. Specifically, our two-step generative lighting estimation pipeline ensures high-quality and robust results under multiple environmental lighting conditions, including various light intensities and color temperatures. Our novel design uses pre-trained large generative models and runtime AR environmental context data to generate HDR environment maps more accurately from limited RGB observations. Our runtime generation control components improve the quality of lighting and reduce estimation latency. Comprehensive quantitative evaluation and user study confirm the accuracy and robustness of CleAR compared to recent generative models.

### v2

@@ 6.3. Perceptual User Study
We conduct an online user study, via Qualtrics [^7] Qualtrics: https://www.qualtrics.com/, to assess the impact of lighting estimation on perceived rendering quality. Our study was approved by our organization’s Institutional Review Board (IRB) and then distributed through personal and professional networks.
6.3.1. Study Protocol
@@
Figure 13. Human perceptual preference comparisons. We compare virtual object rendering quality between CleAR and four other methods. Based on responses from 31 participants, CleAR received the highest average rating of 4.87, outperforming the second-best method, StyleLight, by 12%. CleAR also shows more consistent quality, with a lower standard deviation.
@@
Our survey consists of three chapters. The first chapter surveys the participants’ past experiences with mobile AR and their impressions of virtual object rendering qualities with existing mobile AR applications. The second chapter is a training section that guides participants in completing the quality assessment study. This chapter first shows a quality rating question where the participant will be given five images of the same virtual sphere rendered under lighting conditions from five different sources. Participants are instructed to rate the visual qualities using a Likert scale of 1 to 7, where 1 represents the lowest and 7 represents the highest quality. To avoid bias, participants are not informed which lighting source was used for each image. An example rating for this question is shown to the user for training purposes. Next, we will show a follow-up question to ask participants for feedback, specifically on the quality issues of images generated by our system. The last chapter is the formal quality assessment, consisting of six question groups. Each group uses different environments and virtual object setups, representing environments with varying lighting conditions and objects with various geometries and materials. Six virtual objects are used in total, including a diffuse sphere, a matte sphere, a mirror sphere, a floor lamp, a tea set, and a coffee mug. Specifically, lighting conditions used in these questions were generated using CleAR, DiffusionLight (Phongthawee et al., 2023), StyleLight (Wang et al., 2022), OmniDreamer (Akimoto et al., 2022), and physically captured HDR environment maps. Each group follows the same format used in the training chapter. The lighting conditions used to generate the images are randomized in order and not revealed to participants.
6.3.2. Results and Analysis.
We received responses from 31 participants (48% Respondents are familiar with multiple areas of graphics technology, most commonly image editing (51.6%), followed by video editing (41.9%) and 3D modeling and rendering (35.5%). 25.8% reported no familiarity with specific graphics technologies. 58% had used AR devices, most commonly mobile phones or tablets (48%), followed by Meta Quest (29%).
@@
Figure 13 shows the visual quality assessment results from the last chapter. Excluding ratings for objects rendered with physically captured lighting, participants rated CleAR-generated lighting as producing better renderings for 4 out of the six virtual objects. The visual quality rating of CleAR ranked the top with an average score of 4.87 across all responses. This score is about 12% better than that of the second-best method, StyleLight, which received an average rating of 4.35. Additionally, CleAR has a lower overall standard deviation in ratings (1.56) compared to StyleLight (1.67), indicating more robust estimation quality.
@@
Interestingly, in two question groups—the tea set and coffee mug—participants rated CleAR ’s renderings as higher quality than those produced using physically captured environment maps. Upon further inspection, we found that, while physically captured environment lighting more accurately reflects the radiance conditions in the scene, human perception tends to favor the slightly brighter outputs generated by CleAR. We hope this finding encourages future research to design and evaluate lighting estimation systems with considerations of human perceptual preferences.
@@ 7. Related Work
Environment Lighting and Rendering. Pioneering works have established several ways of capturing high-fidelity physical environment lighting and representing it in digital formats. Omnidirectional HDR environment map is a commonly adopted solution (Debevec, 2006) for representing the environment lighting because it stores the incoming radiance information that can be easily integrated with modern computer graphics rendering. Environment maps can be commonly captured using mirror balls, 360 cameras, or bracketed image stitching (Debevec and Malik, 2023; Brown and Lowe, 2007). A recent work, GLEAM (Prakash et al., 2019), incorporates the mirror ball-assisted environment lighting capturing process into AR applications. However, the requirement for the presence of the mirror ball limits the practicality of AR device usage. Instead, our system uses the generative estimation approach to provide flexible, high-quality lighting estimation. Additionally, HDR environment maps can be challenging to obtain on AR devices due to many sensor limitations. Recent lighting estimation systems (Zhao and Guo, 2021; Zhao et al., 2022a) can only output LDR ones. But our novel two-step generative estimation design allows CleAR to support HDR lighting estimation.
@@
Image Generative Models. Recently, generative models have demonstrated impressive capabilities in image synthesis, as well as generating audio, 3D models, and other data modalities (Yang et al., 2023; Croitoru et al., 2023). The generative diffusion model (Ho et al., 2020) has attracted significant attention from the research community because of its high-quality image generation capabilities. The generative diffusion model uses a physical diffusion process inspired by the Gaussian signal denoising generation process. Combining this novel generation process with the large model parameter sizes, generative diffusion models outperform generative models with prior architectures, such as GAN (Goodfellow et al., 2020) or VAE (Esser et al., 2021). Several recent works (Chen et al., 2022; Akimoto et al., 2022; Wang et al., 2022; Phongthawee et al., 2023) also propose to adopt generative models of different architectures to solve the generic lighting estimation problem. Most notably, DiffusionLight (Phongthawee et al., 2023) achieves state-of-the-art estimation accuracy with diffusion models. In this work, we present a generative model-based approach to lighting estimation tailored for mobile AR applications, addressing key challenges in achieving visually coherent and fast generation.
@@
Context-Aware Mobile System. Context-aware computing is a classical computing paradigm in which applications sense and adapt to contextual information (Chen and Kotz, 2000). Early research in context-aware AR systems (Starner et al., [n. d.]) demonstrated that important environment information can be extracted from camera frames for task planning and decision-making. In recent years, new developments of AR systems have also sought to leverage broad types of environment context information to assist several AR tasks and achieve better user experiences (Lam et al., 2021; Tahara et al., 2020; Wang et al., 2020). Our work leverages four types of AR context data—camera RGB, scene semantic map, and ambient light intensity and color temperature—to guide environment map estimation and align virtual rendering with real-world lighting.
8. Discussion
Application Use Cases. CleAR’s design enables a wide range of mobile AR applications that demand visually coherent environment lighting. For instance, CleAR can significantly enhance virtual object rendering in entertainment applications by ensuring that rendered objects consistently blend with real-world scenes under diverse lighting conditions. Additionally, CleAR is also well suited for commercial applications. For example, in interior retail applications, CleAR enables the creation of visually appealing virtual furniture that accurately reflects realistic ambient lighting. Finally, CleAR’s tight integration with mobile AR can open new opportunities for versatile solutions in digital production, cinematic content creation, and immersive storytelling.
Implications for Privacy and Security. Our work demonstrates the feasibility and effectiveness of generating high-quality environment lighting information from limited camera observations. While this capability can significantly enhance the realism of virtual object rendering, it also introduces potential risks to user spatial privacy, as the reconstructed environment maps may inadvertently expose details about the user’s physical surroundings. For example, recent work demonstrates that sensitive user information can inadvertently be captured by advanced lighting estimation systems and leaked in AR streaming applications (Zhao et al., 2022b). Additional types of information, such as users’ location, may be inferred from the completed LDR environment map (Guzman et al., 2021). Future work should consider defense mechanisms to prohibit the leakage of sensitive user spatial privacy information.
Limitations and Future Directions. While CleAR demonstrates strong performance in generating high-quality environment lighting, it can be further improved in two main ways. First, a useful extension of CleAR would be to support spatially variant lighting estimation, where lighting conditions differ across physical locations. Supporting this capability requires two key changes to the lighting estimation workflow: (i) enabling 3D reconstruction of the surrounding environment to allow spatial transformation and alignment of environment observations, and (ii) training generative lighting models on spatially variant environment map datasets, which often contain diverse distortion patterns. Existing solutions, such as the near-field reconstruction method in LitAR (Zhao et al., 2022a), can be leveraged to develop 3D environment reconstruction methods. However, the primary challenge for training such generative models lies in the lack of high-quality datasets that capture spatially varying lighting. Future research could explore new data curation pipelines or integrate 3D transformation operations directly into generative models. Second, CleAR’s temporal consistency mechanism can be extended beyond the color appearance refinement to better support long AR sessions and scenes with dynamic lighting changes. Future work may investigate methods for progressive environment map updates to enhance consistency and realism across time.
9. Conclusion
@@
In this work, we introduced a lighting estimation framework CleAR that can be easily integrated into many existing mobile AR applications. Our novel two-step generative lighting estimation pipeline ensures high-quality and robust results under multiple environmental lighting conditions, including various light intensities and color temperatures. Our design uses pre-trained large generative models and AR context data to generate HDR environment maps more accurately from limited environment observations. Our real-time refinement steps enhance the quality of lighting and improve responsiveness to estimation. Comprehensive quantitative evaluation and user study confirm the accuracy and robustness of CleAR compared to recent generative models.
Appendix A Lighting Property Measurement Details
We describe additional details for the environmental lighting conditions measurement study described in §3. Our measurement focuses on two properties of environmental lighting conditions: light intensity and color temperature, representing an environment’s overall brightness and color appearances. We choose these two lighting condition properties because of their significant impact on virtual object rendering of all material kinds (Zhao et al., 2022a) We select two standard lighting estimation data sources: the Laval dataset (Gardner et al., 2017), an academic open research dataset, and the PolyHaven website (pol, 2024), a royalty-free HDR environment map data website. In total, we collected 2235 and 196 data items from the Laval dataset and PolyHaven, respectively. Following (Phongthawee et al., 2023; Wang et al., 2022), we perform a standard data preprocessing procedure of a color correction process with gamma correction with $\gamma=2.4$ and setting the 99th percentile of pixel intensity to 0.9.
a Light Intensity
b Color Temperature
Figure 14. Lighting condition measurement. We measure the distributions of light intensity (14a) and color temperature (14b) on environment maps collected from the Laval indoor dataset and PolyHaven and their augmented variants. Compared to the original data, our augmentation technique effectively creates greater diversity in the lighting features of the dataset.
where $i$ represents the i-th pixel in the environment map with $N$ pixels. For color temperature value extraction, we first calculate the average pixel RGB value from each environment map and then calculate its correlated color temperature using a recent method (Ohno, 2014). The calculated environment map image color temperature values are given in Kelvin. In Figure 14, we visualize the measured lighting condition distributions. While both datasets contain a wide selection of lighting conditions, the data distribution shows several biases. Predominantly, neutral lighting conditions of light intensity and color temperatures are seen in two datasets. For light intensity, more low-light environments than bright-lighting conditions data items are seen in the Laval dataset. As for the color temperature distribution, a significant part of the PolyHaven data items lies in the cool color range. These data distribution biases have also led to biased training and evaluation of lighting estimation systems.
We also evaluated the augmented versions of the Laval and PolyHaven datasets generated using our data augmentation technique (§B.3). The results demonstrate that our augmentation method successfully produces new dataset variants with more diverse and evenly distributed lighting characteristics. These augmented datasets are well-suited for assessing the robustness of lighting estimation systems under a wide range of lighting conditions.
Appendix B Dataset Generation Details
Below, we describe how we construct the training datasets for the LDR environment map completion and high-intensity map estimation tasks, from three existing open-source datasets.
B.1. LDR Environment Map Completion Data
Next, we mask the LDR environment map images at random angles to generate AR camera observation images. The masks are generated using the pin-hole camera model to simulate the real-world partial environment observations. We also combine multiple image masks to simulate multi-view environment observations. Specifically, the number of views is randomly chosen from 1 to 5, and the camera’s horizontal FoV is randomly chosen between 60 and 120 degrees. For environment semantics, we use a pre-trained semantic map estimation model (Xiao et al., 2018) to estimate semantic maps directly from the collected LDR environment maps. We also use the LDR image mask to generate masked semantic maps representing the environment’s semantic information received by AR devices. For the ambient lighting condition prompt generation, we create the ambient light property labels using the masked partial environment map image pixel values and our defined system threshold values in §4.2.1.
B.2. High-intensity Map Estimation Data
Our novel learning objective enables the high-intensity map estimation model to leverage its pre-trained knowledge on LDR images, allowing it to be effectively fine-tuned using a limited amount of HDR environment map data. This capability is crucial for training with the small yet high-quality HDR dataset provided by the Laval dataset. Since the original Laval dataset contains HDR environment map images, we applied Equation 1 to convert them into 1,489 paired samples of LDR environment maps and their corresponding high-intensity maps. Specifically, we first clamp the raw HDR environment map to the range $[1,+\infty]$ to isolate pixel intensities beyond the LDR range, and then rescale the values to $[0,+\infty]$ to obtain the high-intensity map $I_{i}$. Next, we convert the range of the raw high-intensity $[0,+\infty]$ to the standard LDR pixel range in $[0,1]$. Note that although this transformation is not lossless, our evaluation suggests minimal impact on the overall lighting estimation. In addition to the generated image pair data, we add the text prompt P2 during training.
B.3. Robustness Evaluation Data Generation
Evaluating the robustness of lighting estimation systems is particularly challenging because it requires controlled lighting condition changes. To avoid costly real-time data capturing, we propose an image editing-based method that creates variants of standard lighting estimation testing datasets to represent diverse lighting conditions while maintaining the original environment map visual context. Specifically, our method includes the following three steps. First, we generate a set of edited variants of the Laval dataset by applying a uniform scaling term $s$ to all environment map images. Then, we measure the total light intensity and average color temperatures using the measurement method introduced in §A. Finally, we uniformly sample from the generated data to ensure equal representation of each edited environment map variant across different light intensity and color temperature ranges. The scaling term $s$ are selected from $[0.25,4]$ using a step of $0.125$. For the light-intensity editing, the scaling term is uniformly applied to all three color channels, while only red and blue channels are scaled with $s$ and $1/s$ for color temperature editing (Afifi and Brown, 2020). Figure 8 shows examples of the edited environment maps. We created a set of augmented Laval indoor datasets using this method, with edited average light intensities and color temperatures in the ranges of $[-20\%,20\%]$ and $[-50\%,50\%]$. In total, we generated approximately 60,000 data items from the 290 original environment maps in the test split of the Laval dataset.
Appendix C Model Training Details
Appendix D On-device Real-time Refinement Details
In this section, we show additional results and technical details of the color appearance refinement technique in CleAR. In the refinement process, our technique seeks to create a color refinement matrix of pixel color multipliers for the completed LDR environment maps. The color refinement matrix consists of two types of multiplier values: (i) the global color multiplier and (ii) the local color multiplier. The global color multiplier adjusts the overall image colors of completed environment map images, while the local color multiplier adjusts fine-grained colors on the observed environment map regions. The two color refinement terms are combined into a color refinement matrix and applied to the completed LDR environment maps.
Figure 16. Examples of environment map refinement over an AR session. We demonstrate the effectiveness of our refinement step using selected frames from an AR session in the ARKitScene (Baruch et al., 2021) dataset. In this example, the initial environment map is estimated by CleAR at time $t_{0}$, and subsequently refined at $t_{1}$ through $t_{4}$ using the corresponding AR camera images. As the AR session progresses, the ambient lighting intensity varies. The color refinement step in CleAR effectively adapts the estimated environment maps to reflect these changing lighting conditions. Note that the original AR camera video has been lightly edited to enhance intensity variations for improved visual clarity.
Our technique calculates the global color multiplier by deriving per-channel multipliers as the ratio between average colors in the estimated and observed environment map images. The local multipliers are calculated by first splitting the estimated environment map images and the partial environment observation image into $NxM$ patches. Then, our technique calculates local multipliers as the average color ratios for each patch. A 3x3 median blur filter is used to address color smoothness between the observation edges. Empirically, we found 8x8 is the best image patch size.
We show the effectiveness of our refinement technique via two visual quality evaluation. As shown in Figure 15, our refinement technique can adjust environment map colors even in very challenging lighting conditions. Furthermore, our design choices of median blurring and patch splitting are useful to smooth out the artifacts on the edges. Figure 16 illustrates the effectiveness of this refinement process over time during an AR session. The example video is sourced from the ARKitScene (Baruch et al., 2021) dataset, with minor edits applied to enhance lighting intensity variations. CleAR first estimates the environment map at frame $t_{0}$, and subsequently refines it at frames $t_{1}$ through $t_{4}$ using the corresponding AR camera images. The refined environment maps clearly reflect the observed changes in ambient lighting intensity. This refinement step is essential for maintaining coherence between the physical environment and the rendered virtual content as lighting conditions change over time. In practice, it is especially effective for adapting to minor lighting changes that do not involve significant alterations to scene geometry or object placement.